Calculate the value of the reaction quotient, Q, of silver iodate (AgIO3) when 10.0 mL of 0.018 M AgNO3 is mixed with 10.0 mL of 0.018 M NaIO3. (Please give your answer with 2 significant figures.)




1
Expert's answer
2021-09-15T02:45:03-0400

10.0 mL of 0.018 M AgNO3 

10.0 mL of 0.018 M NaIO3.

Reactions are:-

AgNO3(aq)+NaIO3(aq)"\to" AgIO3(s)+NaNO3(aq)

AgIO3(s)"\leftrightarrow" Ag+(aq)+IO3-(aq)

Q(Reaction quotient) =[Ag+][IO-3]

Moles of AgNO3=Molarity ×Volume

"\frac{10.0 mL ×0.018 mol/L}{1000mL/L}=" 1.8×10-4moles


Using moles ratio,1:1

AgNO3:NaIO3

Moles of NaIO3=1.8×10-4moles

Total volume after mixing =10mL+10mL=20mL


Concentration of [Ag+]=

"\frac{1000mL/L}{20mL}" ×1.8×10-4mol

=9×10-3M

Using mole ratio:-

[Ag+]:[IO-3]

1:1

Therefore the concentration of [IO-3]

=9×10-3M

Reaction quotient(Q)=[Ag+][IO-3]


=[9×10-3][9×10-3]

=81×10-6

Q=8.1×10-5